OpenStudy (anonymous):

For the function 3x – 2y = 12, what is the x-intercept? -4 4 6 -6

OpenStudy (anonymous):

b

OpenStudy (anonymous):

set Y equal to zero and solvefor X!

OpenStudy (smored):

\[3x-2(0)=12\] \[3x-0=12\] \[3x=12\] \[x=4\]

OpenStudy (tkhunny):

3x – 2y = 12 Divide by 12 \(\dfrac{x}{4} + \dfrac{y}{-6} = 1\) x-intercept (4,0) y-intercept (0,-6) Slope -6/(-4) = 3/2 There is good information in the Intercept Form.